OpenStudy (anonymous):

6 over x+4 = 12 over 5x-13 6 = 12 x+4 5x-13

OpenStudy (hba):

\[\frac{ 6 }{ x+4 }=\frac{ 12 }{ 5x-13 }\] Cross multiply :)

OpenStudy (anonymous):

so 12(x+4) = 6(5x-13)

OpenStudy (hba):

Yeah good Now see.  Multiply them and remember (-)(-)=+ (+)(-)=- (-)(+)=- (+)(+)=+